Lincoln died 142 years ago. Each president of the United States can serve for four to eight years. Assume that all presidents have been in office between four and eight years. How many presidents could have served since the time of Lincoln's death?
Explain the steps you would take to solve this problem.

Answer by richwmiller(17219)

2014-1863=151
151/4=37.75 round up 38 possible presidents since Lincoln based on 151 years
142/4=35.5 round up 36 possible presidents since Lincoln based on 142 years
Obama is #44 but 43rd person to be president
Lincoln was #16
There have been 44-16=28 presidents since Lincoln